Let V be a finite-dimensional vector space over F. For every subset SCV, define Sº = {f eV f(s) = 0 Vs ES}.(a) Prove that sº is a subspace of V* (S may not be a subspace!) (b) If W is a subspace of V and r & W, prove that there exis

just part c,d, and e please!!
Answer #1

oeer in det dim (v=n. and {4, r. set {cht, basis for V a basis of vi => D (21) = { di ri(vi) = L; [by ②1 finite dimentional r4:N V** defined by, y (2) linearity bet ne,, V2 EV and & E WLF. by partic) 4 (V, + & 2) = ud, +Kr2 hare =îg+x. 7 = yle) + 24(e & het, sov. Lele, sep so = { s EU* Ifess=0, & RES}. ft so also let re Spon (S) det = f (20)=0, om tre)=0 & ve span(s) = (s